OpenStudy (kingswank):

MEDALS WILL BE GIVEN!!! Use the net to find the lateral area of the cylinder. R=5in H=9in

OpenStudy (tkhunny):

The lateral surface area is everything but the lids. \(LSA = 2\pi RH\). You are given R and H. You're nearly done.

OpenStudy (kingswank):

@tkhunny I just need some help because I don't know any of this

OpenStudy (tkhunny):

Yes, you do. I do not believe you do not know the substitution principle. You have \(2\pi RH\). You are given R = 5 in. Substitute \(2\pi (5 in)H\). You are given H = 9 in. Substitute \(2\pi (5 in)(9 in)\). You need to be able to do that or you WILL fail.

Elsa213 (elsa213):

Great job @tkhunny

OpenStudy (kingswank):

@tkhunny so is the answer 90?

OpenStudy (tkhunny):

No. It has some relationship to the number 90. Where did the \(\pi\) go? Where did the units go? Where did you get 90?

OpenStudy (kingswank):

So is it 45